# TaQ-DiT: Time-aware Quantization for Diffusion Transformers

### v1

@@ IV-B Comparisons With SOTA Works
| Method | FID | sFID | IS | Precision |
| Full Precision* | 5.00 | 19.02 | 274.78 | 0.8149 |
| PTQ4DM* [15] | 89.78 | 57.20 | 26.02 | 0.2146 |
| Q-Diffusion* [17] | 54.95 | 36.13 | 42.80 | 0.3846 |
| PTQD* [16] | 55.96 | 37.24 | 42.87 | 0.3948 |
| RepQ* [20] | 26.64 | 29.42 | 91.39 | 0.4347 |
| PTQ4DiT* [10] | 7.75 | 22.01 | 190.38 | 0.7292 |
| DiTAS [19] | 6.86 | 19.64 | 218.04 | 0.7638 |
| Q-DiT [18] | 6.40 | 18.60 | 211.72 | 0.7609 |
| Ours | 6.03 | 18.50 | 232.36 | 0.7769 |
@@
As shown in Table III, TaQ-DiT surpasses all baselines, validating our effectiveness. Specifically, (1) when compared with quantization methods tailored for U-Net-based DMs, TaQ-DiT outperforms them by a huge margin, e.g., FID. This highlights the need for DiT-specialized quantization algorithms. Besides, (2) TaQ-DiT also exhibits superior results compared to DiT-dedicated quantization baselines. For example, we can offer and on FID and IS, respectively, when compared to the most competitive baseline Q-DiT [18]. We attribute the improvements to (i) the joint reconstruction method to boost overall performance and (ii) our momentum-based shifting and reconstruction-driven migration to facilitate quantization for sensitive Post-GELU activations.
Besides, as shown in Fig. 5, we achieve significantly higher generative quality than other quantization baselines, further highlighting our superiority from a qualitative perspective.
@@ IV-C Ablation Study
| Joint Reconstruction | Shifting + Migration | |||||
@@
Effectiveness of Proposed Methods. As shown in Table IV, (1) our joint reconstruction method results in significant performance improvements, e.g., FID and IS. (2) Incorporating it with our channel-wise shifting and migration methods further enhances the performance, e.g., FID and IS, which validates our effectiveness.
Comparison With Dynamic Quantization. By integrating momentum-base shifting and reconstruction-driven migration, as seen in Table V, our time-variance-aware static approach achieves comparable performance ( FID and IS) compared to the quantization-inefficient dynamic counterpart, which needs online computations for each denoising step. This highlights our effectiveness in accommodating temporal variability while enhancing quantization efficiency.
@@
| Dynamic | 6.16 | 18.74 | 235.05 | 0.7840 |
| Time-Variance-Aware Static (Ours) | 6.03 | 18.50 | 232.36 | 0.7769 |
The top-2% channels with the largest ranges of shifted activations are regarded to be split or migrated.
@@ V Conclusion
In this paper, we have proposed, developed, and validated TaQ-DiT, an advanced post-training quantization framework for DiTs. First, we propose leveraging joint reconstruction of activations and weights to resolve the non-convergence issue. Then, we develop innovative momentum-based shifting and reconstruction-driven migration techniques to facilitate the effective quantization of sensitive Post-GELU activations. Experimental results validate the effectiveness of our method and show its capability of outperforming state-of-the-art quantization methods and narrowing the gap between quantized and full-precision models.

### v2

@@ IV-B Comparisons With SOTA Works
| Method | FID | sFID | IS | Precision | ||
| Full Precision* | 5.00 | 19.02 | 274.78 | 0.8149 | ||
| All layers | PTQ4DM* [14] | 89.78 | 57.20 | 26.02 | 0.2146 | |
| Q-Diffusion* [8] | 54.95 | 36.13 | 42.80 | 0.3846 | ||
| PTQD* [6] | 55.96 | 37.24 | 42.87 | 0.3948 | ||
| RepQ* [10] | 26.64 | 29.42 | 91.39 | 0.4347 | ||
| PTQ4DiT* [19] | 7.75 | 22.01 | 190.38 | 0.7292 | ||
| Ours | 6.66 | 19.27 | 221.61 | 0.7781 | ||
| MHAs + PFs | DiTAS [3] | 6.86 | 19.64 | 218.04 | 0.7638 | |
| Ours | 6.03 | 18.50 | 232.36 | 0.7769 | ||
@@
As shown in Table III, TaQ-DiT surpasses all baselines, validating our effectiveness. Specifically, with all layers quantized following the setting of PTQ4DiT[19]: (1) when compared with quantization methods tailored for U-Net-based DMs, TaQ-DiT outperforms them by a huge margin, e.g., FID. This highlights the need for DiT-specialized quantization. (2) TaQ-DiT also exhibits superior results compared to DiT-dedicated baseline PTQ4DiT[19] with and on FID and IS, respectively. Notably, when quantizing only MHA and PF layers, our method still maintains superior results ( FID and IS) compared to DiTAS[3] under the same setting. We attribute the improvements to (i) the joint reconstruction to boost overall performance and (ii) our time-variance-aware static transformation to facilitate quantization for sensitive Post-GELU activations. Besides, as shown in Fig. 4, we achieve much better visual quality, further highlighting our superiority from a qualitative perspective.
@@ IV-C Ablation Study
| Joint | Time-Variance-Aware | |||||
| Reconstruction | Static Transformation | |||||
@@
Effectiveness of Proposed Methods. As shown in Table IV, (1) our joint reconstruction method results in significant performance improvements, e.g., FID and IS. (2) Incorporating it with our time-variance-aware static transformation methods further enhances the performance, e.g., FID and IS, which validates our effectiveness.
Comparison With Dynamic Quantization. By integrating momentum-base shifting and reconstruction-driven migration, as seen in Table V, our time-variance-aware static transformation achieves comparable performance ( FID and IS) compared to the quantization-inefficient dynamic counterpart, which needs online computations for each denoising step (dynamic shifting and splitting discussed in Method 1 and Method 2 of Sec. III-B, respectively). This highlights our effectiveness in both performance and efficiency.
@@
| Dynamic Shifting and Splitting (Vanilla) | 6.16 | 18.74 | 235.05 | 0.7840 |
| Time-Variance-Aware Static Transformation (Ours) | 6.03 | 18.50 | 232.36 | 0.7769 |
The top-2% channels with the largest ranges of shifted activations are regarded to be split or migrated. We only quantize MHA and PF layers.
@@ V Conclusion
We have proposed and validated TaQ-DiT, an advanced post-training quantization framework for DiTs. First, we propose leveraging joint reconstruction of activations and weights to resolve the non-convergence issue. Then, we develop innovative momentum-based shifting and reconstruction-driven migration techniques to facilitate the effective quantization of sensitive Post-GELU activations. Experimental results validate the effectiveness of our method and show its capability of outperforming state-of-the-art quantization methods and narrowing the gap between quantized and full-precision models.